CALCULATIONS OF THE H_{3}^{+} ROVIBRATIONAL SPECTRUM TO LEVELS NEAR THE BARRIER TO LINEARITY

\underline{JAMES K. G. WATSON}, Steacie Institute for Molecular Sciences, National Research Council, Ottawa, Ontario, Canada K1A OR6.

On the H_{3}^{+} potential surface^{a} the saddle point for linear configurations lies about 9930 cm^{−1} above the lowest vibrational level. For calculations of vibration-rotation energy levels near or above this energy, it is important that the wavefunctions should have the correct behavior at linear geometries. The present work describes calculations using hyperspherical harmonics^{b} as the angular basis. Previous similar calculations of lower energy levels of H_{3}^{+} have been described by Wolniewicz and Hinze^{c}. Here the line intensities as well as wavenumbers for the 5ν_{2} band will be discussed.

aW. Cencek, J. Rychlewski, R. Jaquet, and W. Kutzelnigg, J. Chem. Phys. 108, 2831–6 (1998)

bL. Wolniewicz, J. Chem. Phys. 90, 371–7 (1989)

cL. Wolniewicz and J. Hinze, J. Chem. Phys. 101, 9817–29 (1994)